### v2

@@ 4.4 Bidimensional noise and oracle provided by an algorithmic method
where $f$, defined by $f(y)\triangleq(y_{2}-y_{1}^{3},y_{1}-y_{3}^{3})$ for all $y\triangleq(y_{1},y_{2},y_{3})\in\mathbb{Y}$, is explicitly known while the blackbox $\varepsilon$ is defined by $\varepsilon(x)\triangleq(\frac{\sin(10\pi(x_{2}-x_{1}^{3}))}{5}+\frac{\sin(6\pi(x_{2}-e^{-x_{1}}+1))}{7}+\frac{\sin(12\pi(x_{1}^{2}+x_{2}^{2})^{\frac{1}{2}})}{11})^{2}$ for all $x\in\mathbb{R}^{2}$. The global minimizer of $\varphi$ is $y^{*}\triangleq(0,0,0)$, with $\varphi(y^{*})=0$. In this context, we define the partition sets of $\mathbb{Y}$ as $\mathbb{Y}(x)\triangleq\{y\in\mathbb{Y}:f(y)=x\}$ for all $x\in\mathbb{X}=X\triangleq\mathbb{R}^{2}$, so $\varepsilon_{|\mathbb{Y}(x)}$ is constant. For all $x\triangleq(x_{1},x_{2})\in\mathbb{X}$, some calculation allow us to rewrite $\mathbb{Y}(x)=\{(t,t^{3}+x_{1},\sqrt[3]{t-x_{2}}),\leavevmode\nobreak\ t\in\mathbb{R}\}$, so the unique minimizer of $\varphi_{|\mathbb{Y}(x)}$ is the solution to the problem
@@
where, for any $y\in\mathbb{Y}(x)$$t_{x}(y)$ denotes the unique $t\in\mathbb{R}$ such that $s_{x}(t)\triangleq(t,t^{3}+x_{1},\sqrt[3]{t-x_{2}})=y$. This solution is hardly tractable. However, the problem admits a smooth exact reformulation given by
$\begin{array}[]{cl}\underset{\begin{array}[]{c}M\geq 0,\leavevmode\nobreak\ T\in\mathbb{R}\end{array}}{\operatorname*{minimize}}&\begin{array}[]{l}\varepsilon(x)+M\end{array}\\[4.30554pt] \mbox{subject to}&\begin{array}[t]{ll}T\in I_{x}^{1}(M)\triangleq\left[-M,M\right],\\ T\in I_{x}^{2}(M)\triangleq\left[\sqrt[3]{-M-x_{1}},\sqrt[3]{M-x_{1}}\right],\\ T\in I_{x}^{3}(M)\triangleq\left[-M^{3}+x_{2},M^{3}+x_{2}\right].\end{array}\end{array}$
@@
Figure 8 shows $\varphi$, Figure 9 draws the partition of $\mathbb{Y}$ and the locus of $\gamma$, and Figure 10 (left) and (center) represent respectively $\varepsilon$ and $M$. In this context, $\Phi$ equals
@@
so $M(x)$ is obtainable via a dichotomic search and the singleton $\{T(x)\}$ follows. In practice, running sufficiently many iterations of the dichotomic search provides an accurate approximation of $M(x)$. We define $M^{0}_{\mathrm{inf}}\triangleq\left\lfloor{M(x)}\right\rfloor$ and $M^{0}_{\mathrm{sup}}\triangleq M^{0}_{\mathrm{inf}}+1$, and $\ell\triangleq 0$, and we iterate the dichotomic search on $[M^{\ell}_{\mathrm{inf}},M^{\ell}_{\mathrm{sup}}]$ while $M^{\ell}_{\mathrm{sup}}-M^{\ell}_{\mathrm{inf}}>2^{-30}$. This allows to define $M(x)\approx\hat{M}(x)\triangleq\frac{1}{2}(M^{\ell}_{\mathrm{inf}}+M^{\ell}_{\mathrm{sup}})$, and then $T(x)\approx\hat{T}(x)\triangleq\frac{1}{2}(\min I_{x}(\hat{M}(x))+\max I_{x}(\hat{M}(x)))$, and finally $\gamma(x)\approx\hat{\gamma}(x)\triangleq s_{x}(\hat{T}(x))$.
A total of eight instances of the cDSM are tested. They actually minimize $\hat{\Phi}\triangleq\varepsilon+\hat{M}$ since we approximate $M\approx\hat{M}$ for tractability. Table 4 shows the convergence of these instances towards $x^{*}$. They all return $\left\lVert{\widetilde{x}^{*}-x^{*}}\right\rVert\leq 9E^{-7}$, hence they provide $\widetilde{y}^{*}\triangleq\hat{\gamma}(\widetilde{x}^{*})$ such that $\left\lVert{\widetilde{y}^{*}-y^{*}}\right\rVert\leq 9E^{-7}$ and $\left\lvert{\varphi(\widetilde{y}^{*})-\varphi(y^{*})}\right\rvert\leq 9E^{-7}$. Five instances even return $\left\lVert{\widetilde{x}^{*}-x^{*}}\right\rVert\leq 5E^{-10}$. These behaviours agree with Theorem 2, even if the problem we actually solve slightly differs from the true problem.
@@
Figure 9: Partition sets of $\mathbb{Y}(x)$ and locus of $\gamma$ in the context of Section 4.4.
Figure 10: functions (left) $\varepsilon$, (center) $M$ and (right) $\Phi$ in Section 4.4. In this section $\Phi=\varepsilon+M$.
$x^{0}$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\leq 5E^{-02}$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\leq 5E^{-05}$ $1^{st}\leavevmode\nobreak\ \hat{x}^{k}\leq 5E^{-08}$ returned $\hat{x}^{k}$
$(-2,2)$ $\hat{x}^{032}=3.52E^{-02}$ $\hat{x}^{110}=4.03E^{-05}$ $\hat{x}^{194}=4.79E^{-08}$ $\hat{x}^{220}=1.10E^{-08}$
@@
Table 4: results for eight instances of Algorithm 1 in the context of Section 4.4. In this table, $\hat{x}^{k}\triangleq\lVert{x^{k}}\rVert_{\infty}$.
5 Numerical gain of performance provided by the POf
@@
In this section, we consider again composite problem as in Section 4, where the POf allows to focus on Reformulation ($\mathbf{P}_{\texttt{ref}}$) with $\dim(\mathbb{X})$ small (equal to the number of inputs of the blackboxes $\sigma$ and $\varepsilon$) instead of Problem ($\mathbf{P}_{\texttt{ini}}$) with $\dim(\mathbb{Y})$ possibly large. We alter the problems from Section 4 so that $\dim(\mathbb{Y})\approx 100$ while $\dim(\mathbb{X})\leq 10$. Let us stress that our choice to consider $\dim(\mathbb{Y})$ at most $100$ stands only for ease of presentation and limited computational effort: $\dim(\mathbb{Y})\approx 100$ is already larger than what many DFO algorithms tackle, and our examples could be adapted to match any $0<\dim(\mathbb{X})\leq\dim(\mathbb{Y})\leq+\infty$ (we even have $\dim(\mathbb{X})\leq 4\ll\dim(\mathbb{Y})=+\infty$ in [10]).
@@
In most examples below, computing $\varphi(y)$$\gamma(x)$ and $\Phi(x)$ is very fast for all $y\in\mathbb{Y}$ and $x\in\mathbb{X}$. This likely differs from real cases, where either $\varphi(y)$ may be computed by an expensive computer program and $\gamma(x)$ may be obtained by solving an optimization subproblem. Thus, to establish a realistic comparison between the solvers, we proceed as follows. For each problem, we consider that the cost to compute $\varphi(y)$ is $1$ unit, identical for all $y\in\mathbb{Y}$, and that the computation of $\Phi(x)$ costs $1+\tau$ units, identical for all $x\in\mathbb{X}$ and where $\tau\geq 0$ is fixed. The value $\tau$ captures the relative cost to evaluate $\gamma$ versus those to evaluate $\varphi$, so the computational cost of $\Phi(x)$ models the additional cost required to first compute $\gamma(x)$ before computing $\varphi(\gamma(x))$. Then, for each method, we track the sequence of all evaluated points. This consists in evaluations of $\varphi$ for solvers tackling Problem ($\mathbf{P}_{\texttt{ini}}$) directly, and of $\Phi$ for our method focusing on Reformulation ($\mathbf{P}_{\texttt{ref}}$) instead. Finally, for each method we plot the graph representing the lowest objective value found within a budget of $c$ units, for all $c\in\mathbb{R}_{+}$.
@@
We arbitrarily choose as baseline two popular DFO solvers, NOMAD [6] and PRIMA [30], solving Problem ($\mathbf{P}_{\texttt{ini}}$). We compare them to our strategy formalized in Theorem 2. This shows that if $\tau$ is not prohibitively large, then our method significantly outperforms both NOMAD and PRIMA. Note that this observation does not discredit these solvers: they are not designed for such large-dimensional contexts, and more suited DFO techniques [19, 21, 22, 23, 26] may also perform poorly when we rise $\dim(\mathbb{Y})$ sufficiently high (up to $\dim(\mathbb{Y})=+\infty$ as in [10]). Similarly, our results do not give credit to our naive implementation of the cDSM. What makes our methodology efficient is the sheer dimension reduction provided by the POf, and the guarantee of (possibly, generalized) local optimality of points returned by any algorithm with a covering step.
@@
We follow the extreme barrier approach [3] that minimizes $\overline{\varphi_{\Omega}}$ over $\mathbb{Y}$ instead of $\varphi$ over $\Omega$, and we provide relevant box constraints to NOMAD and PRIMA. We run NOMAD with default parameters, PRIMA with the BObyQA algorithm (to handle the box constraints), and the cDSM as in Remark 1. We give to NOMAD and PRIMA six random starting points $(y^{0}_{\ell})_{\ell\in\llbracket 1,6\rrbracket}$, and we give to our cDSM the starting points $(\chi(y^{0}_{\ell}))_{\ell\in\llbracket 1,6\rrbracket}$. In this section, we denote by $\mathds{1}$ the vector with all entries equal to $1$.
@@
$\sigma:\left\{\begin{array}[]{ccl}\mathbb{R}&\to&\mathbb{R}^{101}\\ x&\mapsto&\left(\leavevmode\nobreak\ x\leavevmode\nobreak\ ,\leavevmode\nobreak\ \left(2\left(1+\frac{i-1}{5}\right)\left\lfloor{\frac{x}{i}}\right\rceil\right)_{i=1}^{25}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \left(25\sin\left(\frac{i-25}{5}\pi x\right)\right)_{i=26}^{50}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \left(x-\frac{10}{i}\right)_{i=51}^{75}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \left(\frac{i}{10}\right)_{i=76}^{100}\leavevmode\nobreak\ \right).\end{array}\right.$
We consider $\sigma$ and $\varepsilon$ as blackboxes so their expression is assumed to be unknown to all methods, but we remark that the global minimizer of $\varphi$ is $y^{*}\triangleq\sigma(0)$, with $\varphi(y^{*})=0$. We define $\mathbb{Y}(x)\triangleq\{x\}\times\mathbb{R}^{100}$ for all $x\in\mathbb{X}\triangleq\mathbb{R}$, since therefore $\varphi_{|\mathbb{Y}(x)}$ admits an easily tractable global minimizer as
@@
Reformulation ($\mathbf{P}_{\texttt{ref}}$) is a $1$-dimensional DFO problem that focuses on $y_{0}$, and the POf highlights how to handle the $100$ others in the original problem. Even by considering $\tau=100$, the POf provides a sensible gain of performance: our naive cDSM relying on the POf outperforms the solvers solving the original problem. Its returned solution has a significantly lower objective value; and the computational budget is comparable if the pointwise cost to evaluate $\gamma$ is $100$ times greater than those of $\varphi$. Our cDSM converges towards the global minimizer (the returned objective value is around $1E^{-8}$ is all cases), while NOMAD and PRIMA solving the original problem remain both in an exploratory phase that returns poor results. The early stop of PRIMA in the first case even shows that PRIMA interrupted itself on a poor incumbent solution with $40\%$ of its budget remaining.
Figure 11: Comparison of the best solution found by each solver depending on the computational budget spent. The six starting points $y^{0}_{i}$$i\in\llbracket 1,6\rrbracket$, are chosen as $6$ observations of the random uniform independent distribution over $[-30,30]^{101}$.
@@ 5.2 Radial noise in a 101-dimensional problem
The results are similar to those in Section 5.1. Even by considering $\tau=100$, the POf provides a gain of performance. Within a comparable budget, our cDSM converges towards the global minimizer, while NOMAD and PRIMA solving the original problem remain in an exploratory phase that returns irrelevant points. The very early stop of NOMAD in three cases even shows that NOMAD sometimes interrupts itself early on its poor incumbent solution.
Figure 12: Comparison of the best solution found by each solver depending on the computational cost spent. The six starting points $y^{0}_{i}$$i\in\llbracket 1,6\rrbracket$, are chosen as $6$ observations of the random uniform independent distribution over $[0,2\pi]^{101}$.
5.3 Noise affected by a nonlinear combinations of all 100 variables
Consider $\mathbb{Y}=Y\triangleq\mathbb{R}^{100}$, and $\Omega\triangleq\{y\in\mathbb{Y}:\forall i\in\llbracket 1,100\rrbracket,\leavevmode\nobreak\ y_{i+1}\geq y_{i}>0\}$, and

### v3

@@ 4.4 Bi-dimensional noise and oracle provided by an algorithmic method
where $f$, defined by $f(y)\triangleq(y_{2}-y_{1}^{3},y_{1}-y_{3}^{3})$ for all $y\triangleq(y_{1},y_{2},y_{3})\in\mathbb{Y}$, is explicitly known while the blackbox $\varepsilon$ is defined by $\varepsilon(x)\triangleq(\frac{\sin(10\pi(x_{2}-x_{1}^{3}))}{5}+\frac{\sin(6\pi(x_{2}-e^{-x_{1}}+1))}{7}+\frac{\sin(12\pi(x_{1}^{2}+x_{2}^{2})^{\frac{1}{2}})}{11})^{2}$ for all $x\in\mathbb{R}^{2}$. The global minimizer of $\varphi$ is $y^{*}\triangleq(0,0,0)$, with $\varphi(y^{*})=0$. In this context, we define the partition sets of $\mathbb{Y}$ as $\mathbb{Y}(x)\triangleq\{y\in\mathbb{Y}:f(y)=x\}$ for all $x\in\mathbb{X}=X\triangleq\mathbb{R}^{2}$, so $\varepsilon_{|\mathbb{Y}(x)}$ is constant. For all $x\triangleq(x_{1},x_{2})\in\mathbb{X}$, some calculation allows us to rewrite $\mathbb{Y}(x)=\{(t,t^{3}+x_{1},\sqrt[3]{t-x_{2}}),\penalty 10000\ t\in\mathbb{R}\}$, so the unique minimizer of $\varphi_{|\mathbb{Y}(x)}$ is the solution to the problem
@@
where, for any $y\in\mathbb{Y}(x)$, $t_{x}(y)$ denotes the unique $t\in\mathbb{R}$ such that $s_{x}(t)\triangleq(t,t^{3}+x_{1},\sqrt[3]{t-x_{2}})=y$. This solution is hardly tractable. However, the problem admits a smooth exact reformulation given by
$\begin{array}[]{cl}\underset{\begin{array}[]{c}M\geq 0,\penalty 10000\ T\in\mathbb{R}\end{array}}{\operatorname*{minimize}}&\begin{array}[]{l}\varepsilon(x)+M\end{array}\\[4.30554pt] \mbox{subject to}&\begin{array}[t]{ll}T\in I_{x}^{1}(M)\triangleq\left[-M,M\right],\\ T\in I_{x}^{2}(M)\triangleq\left[\sqrt[3]{-M-x_{1}},\sqrt[3]{M-x_{1}}\right],\\ T\in I_{x}^{3}(M)\triangleq\left[-M^{3}+x_{2},M^{3}+x_{2}\right].\end{array}\end{array}$
@@
Figure 8 shows $\varphi$, Figure 9 draws the partition of $\mathbb{Y}$ and the locus of $\gamma$, and Figure 10 (left) and (centre) represent respectively $\varepsilon$ and $M$. In this context, $\Phi$ equals
@@
so $M(x)$ is obtainable via a dichotomic search and the singleton $\{T(x)\}$ follows. In practice, running sufficiently many iterations of the dichotomic search provides an accurate approximation of $M(x)$. We define $M^{0}_{\mathrm{inf}}\triangleq\left\lfloor{M(x)}\right\rfloor$ and $M^{0}_{\mathrm{sup}}\triangleq M^{0}_{\mathrm{inf}}+1$, and $\ell\triangleq 0$, and we iterate the dichotomic search on $M^{\ell}_{\mathrm{inf}},M^{\ell}_{\mathrm{sup}}]$ while $M^{\ell}_{\mathrm{sup}}-M^{\ell}_{\mathrm{inf}}>2^{-30}$. This allows to define $M(x)\approx\hat{M}(x)\triangleq\frac{1}{2}(M^{\ell}_{\mathrm{inf}}+M^{\ell}_{\mathrm{sup}})$, and then $T(x)\approx\hat{T}(x)\triangleq\frac{1}{2}(\min I_{x}(\hat{M}(x))+\max I_{x}(\hat{M}(x)))$, and finally $\gamma(x)\approx\hat{\gamma}(x)\triangleq s_{x}(\hat{T}(x))$.
[…]
@@
Figure 9: Partition sets of $\mathbb{Y}(x)$ and locus of $\gamma$ in the context of Section $4.$
Figure 10: Functions (left) $\varepsilon$, (centre) $M$ and (right) $\Phi$ in Section $4.$ In this section $\Phi=\varepsilon+M$.
$x^{0}$ $1^{st}\penalty 10000\ \hat{x}^{k}\leq 5E^{-02}$ $1^{st}\penalty 10000\ \hat{x}^{k}\leq 5E^{-05}$ $1^{st}\penalty 10000\ \hat{x}^{k}\leq 5E^{-08}$ returned $\hat{x}^{k}$
(2,2)(-2,2) $\hat{x}^{032}=3.52E^{-02}$ $\hat{x}^{110}=4.03E^{-05}$ $\hat{x}^{194}=4.79E^{-08}$ $\hat{x}^{220}=1.10E^{-08}$
@@
Table 4: Results for eight instances of Algorithm 2 in the context of Section $4.$ In this table, $\hat{x}^{k}\triangleq\lVert{x^{k}}\rVert_{\infty}$.
5 Gain of performance provided by the DFPOm
@@
In this section, we tackle composite greybox problems where the POf allows for Reformulation ($\mathbf{P}_{\texttt{r}}$) with $\dim(\mathbb{X})$ small (equal to the number of inputs of the blackboxes $\sigma$ and $\varepsilon$) even though $\dim(\mathbb{Y})$ is possibly large. We show that the DFPOm is efficient in this context. We alter the problems from Section 4 so that $\dim(\mathbb{Y})\approx 100$ while $\dim(\mathbb{X})\leq 10$. Let us stress that our choice to consider $\dim(\mathbb{Y})\approx 100$ at most stands only for limited computational effort: $\dim(\mathbb{Y})\approx 100$ is already larger than what many DFO algorithms tackle, and our examples could be adapted to match any $0<\dim(\mathbb{X})\leq\dim(\mathbb{Y})\leq+\infty$ (we even have $\dim(\mathbb{X})\leq 4\ll\dim(\mathbb{Y})=+\infty$ in [11]).
@@
In most examples below, computing $\varphi(y)$, $\gamma(x)$ and $\Phi(x)$ is very fast for all $y\in\mathbb{Y}$ and $x\in\mathbb{X}$. This likely differs from real cases, where either $\varphi(y)$ may be computed by an expensive computer program and $\gamma(x)$ may be obtained by numerically solving an optimization subproblem. Thus, to establish a realistic comparison between the solvers, we proceed as follows. For each problem, we consider that the cost to compute $\varphi(y)$ is $1$ unit, identical for all $y\in\mathbb{Y}$, and that the computation of $\Phi(x)$ costs $1+\tau$ units, identical for all $x\in\mathbb{X}$ and where $\tau\geq 0$ is fixed. The value $\tau$ captures the relative cost to evaluate $\gamma$ versus those to evaluate $\varphi$, so the computational cost of $\Phi(x)$ models the additional cost required to first compute $\gamma(x)$ before computing $\varphi(\gamma(x))$. Then, for each method, we track the sequence of all evaluated points. This consists in evaluations of $\varphi$ for solvers tackling Problem ($\mathbf{P}_{\texttt{i}}$) directly, and of $\Phi$ for our method focusing on Reformulation ($\mathbf{P}_{\texttt{r}}$) instead. Finally, for each method we plot the graph representing the lowest objective value found within a budget of $c$ units, for all $c\in\mathbb{R}_{+}$.
@@
We compare the DFPOm against the DFO solvers NOMAD [7] and PRIMA [35] solving Problem ($\mathbf{P}_{\texttt{i}}$) directly. We observe that the DFPOm significantly outperforms both NOMAD and PRIMA, unless $\tau$ is prohibitively large. Note that this does not discredit these solvers: as we discuss in Section 2, they are not designed for large-dimensional contexts, and others DFO techniques [19, 21, 22, 23, 26] may also perform poorly in contexts suited for the POf. Similarly, our results do shred light on our instance of the cDSM. What makes the DFPOm efficient is the sheer dimension reduction provided by the POf.
@@
We follow the extreme barrier approach [4] that minimizes $\overline{\varphi_{\Omega}}$ over $\mathbb{Y}$, with relevant box constraints. We run NOMAD with default parameters, PRIMA with the BObyQA algorithm, and the cDSM as in Remark 5. We give to NOMAD and PRIMA six random starting points $(y^{0}_{\ell})_{\ell\in\llbracket 1,6\rrbracket}$, and we give to our cDSM the starting points $(\chi(y^{0}_{\ell}))_{\ell\in\llbracket 1,6\rrbracket}$. In this section, we denote by $\mathds{1}\triangleq(1,\dots,1)$.
@@
$\sigma:\left\{\begin{array}[]{ccl}\mathbb{R}&\to&\mathbb{R}^{101}\\ x&\mapsto&\left(\penalty 10000\ x\penalty 10000\ ,\penalty 10000\ \left(2\left(1+\frac{i-1}{5}\right)\left\lfloor{\frac{x}{i}}\right\rceil\right)_{i=1}^{25}\penalty 10000\ ,\penalty 10000\ \left(25\sin\left(\frac{i-25}{5}\pi x\right)\right)_{i=26}^{50}\penalty 10000\ ,\penalty 10000\ \left(x-\frac{10}{i}\right)_{i=51}^{75}\penalty 10000\ ,\penalty 10000\ \left(\frac{i}{10}\right)_{i=76}^{100}\penalty 10000\ \right).\end{array}\right$.
We consider $\sigma$ and $\varepsilon$ as blackboxes, so their expression is assumed to be unknown to all methods, but we remark that the global minimizer of $\varphi$ is $y^{*}\triangleq\sigma(0)$, with $\varphi(y^{*})=0$. We define $\mathbb{Y}(x)\triangleq\{x\}\times\mathbb{R}^{100}$ for all $x\in\mathbb{X}\triangleq\mathbb{R}$, since therefore $\varphi_{|\mathbb{Y}(x)}$ admits an easily tractable global minimizer as
@@
The DFPOm outperforms NOMAD and PRIMA tackling Problem ($\mathbf{P}_{\texttt{i}}$) directly. Our cDSM closely approaches the global minimizer while NOMAD and PRIMA remain in an exploratory phase that returns poor results, and the computational budget is comparable when $\tau=10$.
Figure 11: Comparison of the best solution found by each solver depending on the computational budget spent. The six starting points $y^{0}_{i}$, $i\in\llbracket 1,6\rrbracket$, are chosen as 66 observations of the random uniform independent distribution over $30,30]^{101}$.
@@ 5.2 Radial noise in a 101-dimensional problem
The results are similar to those in Section 5.1. Even by considering $\tau=100$, the POf provides a gain of performance. Within a comparable budget, the cDSM in our DFPOm converges towards the global minimizer, while NOMAD and PRIMA solving the original problem remain in an exploratory phase that returns irrelevant points. The very early stop of NOMAD in three cases even shows that NOMAD sometimes interrupts itself early on its poor incumbent solution.
Figure 12: Comparison of the best solution found by each solver depending on the computational cost spent. The six starting points $y^{0}_{i}$, $i\in\llbracket 1,6\rrbracket$, are chosen as 66 observations of the random uniform independent distribution over $0,2\pi]^{101}$.
5.3 Noise affected by a nonlinear combination of all 100 variables
Consider $\mathbb{Y}=Y\triangleq\mathbb{R}^{100}$, and $\Omega\triangleq\{y\in\mathbb{Y}:\forall i\in\llbracket 1,100\rrbracket,\penalty 10000\ y_{i+1}\geq y_{i}>0\}$, and